OpenStudy (anonymous):

How do I solve this quadratic and linear equation using substitution? I know that the answer is (6,-2) (-9,-47) But I cannot figure out the steps to get the answers. y=3x-20 y=x^2+34 Thank you for your help.

OpenStudy (anonymous):

write \[3x-20=x^2+34\]and solve for \(x\)

OpenStudy (anonymous):

\[ y = 3x - 20 \] substitute this y in the second equation: \[ 3x - 20 = x^2 + 34 \] \[ x^2 + 34 - 3x + 20 = 0 \] \[ x^2 - 3x + 54 = 0 \] \[ x = \frac{3 \pm \sqrt{(-3)^2 - 4 \cdot 1 \cdot 54}}{2} \] Solve for x and you have your answer

OpenStudy (anonymous):

How do you go from the third equation to the fourth one?

OpenStudy (anonymous):

That is quadratic formula. An equation of form \[ ax^2 + bx + c = 0 \] has its roots as: \[ x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} \]

OpenStudy (anonymous):

since you have the solution, you know it factors

OpenStudy (anonymous):

i.e. you already know \(x=-9\) or \(x=6\) so it must factor as \[x^2-3x+54=(x-9)(x+6)=0\]

OpenStudy (phi):

Apparently there is a typo. Instead of y=3x-20 y=x^2+34 use y= 3x-20 y= -x^2+34 (change to -x^2 rather than +x^2) now you should get the answers you posted.